Complete each of the following equations by writing the correct formulas of the products and balancing the equation:
Express your answer as a chemical equation. Identify all of the phases in your answer.

combustion: C4H6(g)+O2(g)⟶Δ ________ + _________
1
Expert's answer
2015-10-13T11:17:08-0400
Combustion of propyne:
C4H6(g) + O2(g) = CO2(g) + H2O(g)
Balanced reaction equation:
2C4H6(g) + 11O2(g) = 8CO2(g) + 6H2O(g)